Menu icon Customer tools: 
System Status

Current System Status

E-Mail Delivery to BT and Yahoo

14:04 17/02/2017

Identified - We are seeing that a number of rejections to BT and Yahoo are being caused by badly formed SPF records. If you are having issues sending to addresses with BT and Yahoo from our systems we recommend using the following SPF:-

v=spf1 exists:%{i}.mta.spf.extendcp.co.uk exists:%{i}.web.spf.extendcp.co.uk -all

If you need any assistance in updating this our support team will be happy to help.
Feb 10, 17:52 GMT

Investigating - We are seeing some failures as well as delays with Mail Delivery to BT and Yahoo addresses from messages sent from our network.

Our System Administrators are currently working on this and we will provide further updates here as we have them.

We apologise for the inconvenience caused as a result of this
Jan 23, 12:30 GMT

Mail 114 - Abnormal server performance

11:21 17/02/2017

Performance on this server has improved, we are still monitoring this and will continue to do so over the weekend to ensure stability.

This status page will be kept up to date with the latest information on this server.

Update - 21/02/2017 - 13:00

Work is ongoing to improve performance and stability, our System Administration Team are working on this but there may still be some degraded performance at this time.

We apologise for the inconvenience caused as a result of this problem.

Update - 21/02/2017 - 15:25

Currently, users will not be able to login to mailboxes hosted on this server. This is intentional whilst work is being carried out on the server - it will not impact the receiving of emails.

Update - 21/02/2017 - 17:15

Users are currently unable to login to Mailboxes on this server while work is being carried out by our System Administration Team. There is a large mail queue on this server and it is still accepting messages for delivery which should be available once logins are restored.

Update - 21/02/2017 - 20:32

Users are still unable to login to Mailboxes on this server while work is being carried out by our System Administration Team. The mail queue has been reduced significantly on this server however further work is still being carried out, this sever is still accepting messages for delivery which should be available once logins are restored. We apologise for any inconvenience caused.

Update - 22/02/2017 - 08:54

Work has continued through the night and is still being carried out by our System Administration Team, however it is in it's final stages. The mail queue has been reduced completely on this server, and a migration of a large amount of mailboxes is currently being finished.

Once completed mail login should be possible again. This sever is still accepting messages for delivery which should be available once logins are restored. We apologise for any inconvenience caused.

Update - 22/02/2017 - 10:25

Migration away of a large number of mailboxes is almost finished and should be complete in the next hour, after this we'll be trailing enabling normal access again.

Update - 22/02/2017 - 11:47

Logging in to mailboxes should now be possible however we've had to limit the number of concurrent connections to prevent the server being placed under a high load. This may result in your login being denied on some attempts, please do not simply keep retrying, give it some time between attempts.

As the server settles we'll increase the number of possible connections and we're continuing to live migrate mailboxes to new servers, this should result in no further interruption to service.

Update - 22/02/2017 - 13:55

The limit to the number of concurrent connections has been lifted and the server is coping well, we hope to be able to lift this again soon.

Update - 22/02/2017 - 15:10

There are still some connection limits in place to keep the server at a safe load, but we are increasing these on an ongoing basis. If you are not able to connect you should be able to after waiting for a short period of time.

Update - 22/02/2017 - 16:20

We are increasing the number of available IMAP and POP connections further to allow more users to login to Mailboxes on this server. Migrations of Mailboxes that are being performed to other servers is underway and is around 50% complete. As these migrations continue we will be able to increase the number of available connections for IMAP and POP.

Update - 22/02/2017 - 19:00

We have further increased the available connections to Mailboxes on this server back to near capacity and we expect the ongoing migration to complete within the next 2 hours.

Customers should now be able to connect to Mailboxes with much better success than earlier in the day.

Update - 22/02/2017 - 22:28

The migrations is taking longer than expected but is still currently running. Overall health of the server is looking better and we will continue to monitor this over the course of the evening/night.

Update - 23/02/2017 - 00:05

The 'load' on the mailserver looks good at present, and looking at the logs customers are able to successfully login to mailboxes. Some may experience timeouts but should be fine after a few retries. The mail queue is at normal or expected levels.

Update - 23/02/2017 - 03:49

Migration work has finished for now, which will mean less load on the mailserver. We may resume work later this morning though, however this will be evaluated first by our senior system administration team.

web62, web63, web64, web65, web71, web72, web73, web74, web193, web194, web195, web196 abnormal server performance

09:39 20/02/2017

We are currently having adverse performance on these servers.

Upon investigating the causes we've taken the decision to migrate a number of the packages on these clusters over to a new database server.

We expect to have it back to normal operation soon, and apologise for any inconvenience caused.

Update - 21/02/2017 - 13:00

Our System Administration Team is continuing to work on improving performance on these servers. While this work is ongoing there may be some periods of slowness accessing some sites.

We apologise for the inconvenience caused as a result of this.

Update - 22/02/2017 - 18:00

Some users may notice poor performance at peak times when working with Database driven websites. Our System Administration Team are aware of this and we are continuing to work on performance improvements.

Update - 23/02/2017 - 07:17

This is still being investigated and worked on as a priority by our system administrator team, we will hopefully have another update later this morning on this issue.

Planned System Maintenance

There is currently no planned maintenance.

© Copyright and legal notices